OpenStudy (anonymous):

Solve for y in the following system of equations. 10x + 2y = 0 -10x - 8y = -60

OpenStudy (anonymous):

Once again: 10x + 2y = 0 + -10x + (-8y) = -60 -------------------- 0x - 6y = -60 -6y = -60 y = 10

OpenStudy (anonymous):

10x +2y = 0 -10x-8y=-60 ----------- -6y=-60 y = -60/-6 = 10 10x +2(10) = 0 10x +20 = 0 10x = -20 x = -20/10 = .-2